OpenStudy (anonymous):

Find the original price of a pair of shoes if the sale price is $48 after a 20% discount......what is the original price!

OpenStudy (anonymous):

when you do a 20% discount what you're left with is 80% of the original price. to get 80% of something you multiply by 0.8 so if our original price was X then our discount price is 0.8X so now 48 = 0.8X divinding by 0.8 we get X = 60

OpenStudy (anonymous):

does that make sense to you?

OpenStudy (anonymous):

if it doesnt just say and ill try to help :)

OpenStudy (anonymous):

yes would all the problems be able to be worked out like that

OpenStudy (anonymous):

all problems of that form :D if you have one that you're unsure about post it and ill have a look

OpenStudy (anonymous):

the number of different cars sold in some country rose from 203 to 1997 to 208 in 2007. find the percent increase round to the nearest whole percent

OpenStudy (anonymous):

this one is slightly different so percentage increase is: \[\frac{\text{change}}{\text{original value}} \times 100\] have a go and if you get stuck ill help :)

OpenStudy (anonymous):

102?

OpenStudy (anonymous):

208/203*100??

OpenStudy (anonymous):

ah you've made a mistake everyone tends to make, the change is not 208, but 208 - 203 = 5

OpenStudy (anonymous):

ok so 208-203=5 so it would be five percent?

OpenStudy (anonymous):

.05*100?

OpenStudy (anonymous):

ahh this seems so simple and i cant fgure it out

OpenStudy (anonymous):

dont worry, with practice it gets easier :) \[\frac{208 - 203}{203} \times 100 = 0.02463054187 \times 100 = 2.46 %\]

OpenStudy (anonymous):

see there, the change is 5, divided by the original value which was 203

OpenStudy (anonymous):

ok i see let me try one and ill get back to yu

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

find the last years salary if, after a 6% pay raise, this years salary is 33920

OpenStudy (anonymous):

i got the other... :) thanks

OpenStudy (anonymous):

this ones the same situation as the first, in a way. if we increase something by 6% then we will have 106% to actually increase we have to multiply by (106)/(100) = 1.06 lets say the starting salary was Y we then increase it by 6% and we are left with 1.06Y now 33920 = 1.06Y dividing by 1.06 gets us 32000 = Y

OpenStudy (anonymous):

gotcha give me a sec doign another one

OpenStudy (anonymous):

ok got that but heres another i have tried three times

OpenStudy (anonymous):

a coffee company wants a new flavor of cajun coffee. how many pounds of coffee worth $9 a pound should be added to 10 pounds of coffee worth $3 a pound to get a mixture worth $7 a pound

OpenStudy (anonymous):

how wouldi set it up

OpenStudy (anonymous):

so this one its worth making an equation. cost/amount = 7 lets call the amount of pounds we add "x" starting cost is 10*3 = $30 and each time we add a pound it adds $9 so if we add x pounds the value is 30 + 9x starting amount is 10 pounds and it adds 1 pound each time so if we add x pounds then the amount is 10 + x we know that cost/amount = 7 (30 + 9x)/(10 + x) = 7 now you just have to solve this

OpenStudy (anonymous):

39x/10x=7?? would be the next step

OpenStudy (anonymous):

sorry 11x

OpenStudy (anonymous):

no, you're adding the like terms wrong 30 + 9x isnt 39x what you have to do is multiply by the bottom expression: